OpenStudy (anonymous):

Could someone help with this chem question? I did it and got 25 ml, but i feel like its wrong.

OpenStudy (anonymous):

It says to find the volume of 0.20 M NaOH required to adjust the pH of the 0.10 M acetic acid solution to a pka value of acetic acid. I have 50 ml of acetic acid

OpenStudy (abb0t):

Can you show me how you got 25 mL?

OpenStudy (anonymous):

sure @abb0t

OpenStudy (anonymous):

So I know pka =4.76= pH and mole of acetic acid =0.005 mol. The equation become HA + NaOH = NaA +H2O I have HA=0.005 and set NaOH= x and NaA= 0.005-x

OpenStudy (anonymous):

\[pH= pka+ \log \frac{ A }{ HA }\] \[4.76 = 4.76 + \log \frac{ 0.005-x }{ 0.005 }\]\[1 = \frac{ 0.005-x }{ 0.005 }\]

OpenStudy (anonymous):

i did it wrong the first time...and now im stuck